OpenStudy (anonymous):

The diameter of a circle is tripled. How does this affect the area of the circle

OpenStudy (tkhunny):

Just do it. \(Area = \pi r^{2} = \pi \left(\dfrac{d}{2}\right)^{2}\), right? Now, in place of "\(d\)", put "\(3d\)" and see what happens.

OpenStudy (anonymous):

Ok I'm in sixth grade can u make it easier please

OpenStudy (anonymous):

Please

OpenStudy (tkhunny):

No. That's all there is to it. Practice problems: \(3^{2} =\;??\) \((2f)^{2} =\;??\)

OpenStudy (tkhunny):

It doesn't matter what grade you are in. If you should not be given such a problem, why were you given this problem? You must have been given SOME tools to solve it. Did you talk in class about length, area, and volume and their relationships? We have to find some way to figure out why it was reasonable for you to have this problem statement to solve.

OpenStudy (anonymous):

Ik that pi =3.14 But like the question doesn't say what the original area is to the diameter or the radius It just says: The diameter of a circle is tripled. How does this affect the area of the circle?

OpenStudy (anonymous):

Plz help me My mom is terrible at math!!!!

OpenStudy (johnweldon1993):

@Catierose02 Alright so we have a circle... The area of a circle is equal to \(\large \pi r^2\) r is equal to the radius of the circle...which is half of the diameter  

OpenStudy (johnweldon1993):

So if the area if \(\large \pi r^2\) Lets choose a random Diameter....lets say 10 this means that the radius would be half that...and will be 5 Area of that will be \[\large \pi 5^2 = 25 \pi\] Okay....now say that we triple that same diameter....so instead of 10...we have 3 times 10 which is 30 That means the radius will now be half of 30....which is 15 Area of this new one \[\large \pi 15^2 = 225 \pi\] So we went from \(\large 25\pi \) to \(\large 225 \pi\) So what change is that?

OpenStudy (anonymous):

Ummmmm 200 squared units

OpenStudy (johnweldon1993):

Not quite...we are thinking "How many times bigger is this new circle compared to the old circle? So what is 225 / 25 ?

OpenStudy (anonymous):

9

OpenStudy (johnweldon1993):

Correct... so If the diameter of a circle is tripled...how will it affect the area?" Well it would make the area 9 times bigger

OpenStudy (anonymous):

Ooooooooooooooo OMG THANK U SOOOOOO MUCH!

OpenStudy (johnweldon1993):

No problem :)